# InjecGuard: Benchmarking and Mitigating Over-defense in Prompt Injection Guardrail Models

### v2

@@ Abstract
Prompt injection attacks pose a critical threat to large language models (LLMs), enabling goal hijacking and data leakage. Prompt guard models, though effective in defense, suffer from over-defense—falsely flagging benign inputs as malicious due to trigger word bias. To address this issue, we introduce NotInject, an evaluation dataset that systematically measures over-defense across various prompt guard models. NotInject contains 339 benign samples enriched with trigger words common in prompt injection attacks, enabling fine-grained evaluation. Our results show that state-of-the-art models suffer from over-defense issues, with accuracy dropping close to random guessing levels (60%). To mitigate this, we propose InjecGuard, a novel prompt guard model that incorporates a new training strategy, Mitigating Over-defense for Free (MOF), which significantly reduces the bias on trigger words. InjecGuard demonstrates state-of-the-art performance on diverse benchmarks including NotInject, surpassing the existing best model by 30.8%, offering a robust and open-source solution for detecting prompt injection attacks. The code and datasets are released at https://github.com/SaFoLab-WISC/InjecGuard.
@@
Hao Li1,2thanks: These authors contribute equally to this work., Xiaogeng Liu3, 1University of Electronic Science and Technology of China, 2National University of Singapore, 3University of Wisconsin-Madison
@@ 1 Introduction
Prompt injection attacks Perez and Ribeiro (2022); Greshake et al. (2023); Liu et al. (2024) represent a serious and emerging threat to the security and integrity of large language models (LLMs) Brown et al. (2020). These attacks exploit the models’ reliance on natural language inputs by inserting malicious or manipulative prompts, leading to undesirable behaviors such as goal hijacking or sensitive data leakage. For instance, a well-known prompt injection technique involves instructing the LLM to “ignore previous instructions”, Branch et al. (2022); Harang (2023a); Perez and Ribeiro (2022); Willison (2022) which can override built-in safeguards and enable the execution of unauthorized actions. As the deployment of LLMs continues to expand across industries, the risk posed by prompt injection attacks becomes more pronounced, demanding effective and efficient defense mechanisms.
@@
To address it, prompt guard models Meta (2024); ProtectAI.com (2024); Deepset (2024b); fmops (2024); LakeraAI (2024a) have recently been proposed as a promising solution. These models work by analyzing the semantic meaning of the input data to detect malicious intent before it reaches the LLM. Unlike LLM models, prompt guard models are lightweight and computationally efficient, as they do not require the high inference cost associated with LLMs. Additionally, these models operate independently, without the need for victim LLM’s responses, further reducing computation costs compared to approaches like LLM guardrails Inan et al. (2023). These attributes make prompt guard models adaptable to various environments, and an attractive solution in scenarios where speed and resource optimization are critical. We
Figure 2: Left: PromptGuard Meta (2024) 1, Right: ProtectAIv2 ProtectAI.com (2024) 2, which is current SotA prompt guard model. We found that existing prompt guard models have an over-defense issue, i.e., misclassify benign inputs as malicious due to an overreliance on specific trigger words, such as “ignore”.
@@
11footnotetext: https://huggingface.co/meta-llama/Prompt-Guard-86M22footnotetext: https://huggingface.co/protectai/deberta-v3-base-prompt-injection-v2
@@ 2 Related Works
Prompt injection attacks. Prompt injection attacks have become a critical threat to LLMs. Since LLMs process inputs in natural language, they often face challenges distinguishing between legitimate user commands and external manipulative inputs, leaving them vulnerable. The concept is first identified in research by Perez and Ribeiro (2022), revealing that LLMs could be misled by simple, crafted inputs, resulting in goal hijacking and prompt leakage. Several studies have been proposed Greshake et al. (2023); Wang et al. (2023); Pedro et al. (2023); Yan et al. (2023); Yu et al. (2023); Salem et al. (2023); Yip et al. (2024); Zhan et al. (2024); Liu et al. (2024); Pasquini et al. (2024); Shi et al. (2024), addressing various aspects of prompt injection attacks, such as handcrafted methods Toyer et al. (2023), automatic attack algorithms Liu et al. (2024), and benchmarks Liu et al. (2023); Debenedetti et al. (2024). Public discussions Harang (2023b); Willison (2022, 2023) have also underscored the risks of prompt injection attacks on commercial LLMs. Prompt Injection datasets are also introduced, such as PINT LakeraAI (2024b), Safeguard-Injection Erdogan et al. (2024), TaskTracker Abdelnabi et al. (2024), and BIPIA Yi et al. (2023), etc.
@@
Prompt guard models. Prompt guard models offer a promising approach to defend against prompt injection attacks by detecting malicious intent in inputs. These methods are computationally efficient and, unlike LLM guardrails Inan et al. (2023), do not require an additional round of victim LLM inference to generate a response. Several prompt guard models have been proposed recently. Fmops fmops (2024) is a guardrail trained by FmopsAI, utilizing DistilBERT Sanh et al. (2020) as the backbone. Deepset Deepset (2024b), PromptGuard Meta (2024), and ProtectAIv2 ProtectAI.com (2024) all employ DeBERTaV3-base He et al. (2023) as the backbone. Deepset Deepset (2024b) is trained on Deepset-Injection Deepset (2024a), a small open-source dataset with 546 samples. PromptGuard is a model trained by Meta with a closed-source dataset. ProtectAIv2 is also trained with a closed-source dataset by Protect AI, achieving SotA performance among open-source guard models Deepset (2024b); fmops (2024); Meta (2024); ProtectAI.com (2024) on PINT LakeraAI (2024b). Furthermore, LakeraGuard LakeraAI (2024a) is a commercial guard model developed by LakeraAI, with undisclosed model details, such as scale and architecture, and available solely through an API.
@@
However, most of these prompt guard models fmops (2024); Deepset (2024b); Meta (2024); ProtectAI.com (2024) suffer from over-defense issues, where they rely on shortcuts triggered by certain keywords to make predictions, leading them to incorrectly categorize a benign input that contains the keywords as malicious. Moreover, most of these models fmops (2024); Meta (2024); ProtectAI.com (2024); LakeraAI (2024a) are trained using closed-source data with undisclosed implementation or training details. In this paper, we introduce an over-defense dataset to systematically evaluate the over-defense issue, and propose a novel and completely open-source training approach that achieves SotA performance across diverse benchmarks, including the over-defense dataset.
@@ 3 Over-defense Dataset: NotInject
Figure 4: The pipeline for constructing NotInject dataset.
@@ 3.1 The Over-defense Issue
While prompt guards offer several advantages, such as low overhead as previously mentioned, we have identified a critical limitation: they exhibit severe over-defense issues, even in some advanced models Meta (2024); ProtectAI.com (2024). Specifically, these models learn a shortcut from certain trigger words, like “ignore”, directly to the final prediction. As shown in Fig. 3 with red lines, ProtectAIv2 ProtectAI.com (2024), one of the SotA prompt guard models on PINT benchmark, distributes excessive and unbalanced attention to the word “ignore”, leading it to incorrectly categorize a benign input as malicious. To further illustrate, we input a benign sentence containing the word “ignore” into two advanced prompt guard models: PromptGuard Meta (2024) and ProtectAIv2 ProtectAI.com (2024), and present the results in Fig. 2. Both models incorrectly classify the benign instruction containing the word “ignore” as malicious.
@@ 3.2 Construction of NotInject
Algorithm 1 Trigger Words Identification
1:Benign dataset $D_{b}$, Malicious dataset $D_{m}$, Integer $k$
2:Compute word frequencies in $D_{b}$ to get frequency list $F_{b}$
3:Compute word frequencies in $D_{m}$ to get frequency list $F_{m}$
4:Sort $F_{b}$ in descending order to get rank list $R_{b}$
5:Sort $F_{m}$ in descending order to get rank list $R_{m}$
6:for all words $w$ in $R_{b}\cup R_{m}$ do
7:     Compute rank difference $(\Delta r(w)=R_{b}(w)-R_{m}(w))$
8:     Add $(w,\Delta r(w))$ to list $L$
9:end for
10:Sort list $L$ in descending order based on $\Delta f(w)$
11:return Top $k$ words from list $L$
@@
Trigger Words Refinement. Recognizing that automated methods may include irrelevant or common words not indicative of prompt injection attempts, we proceed with a word rectification process. We first employ LLM to automatically filter the list of candidate words to remove any that may not be pertinent. Specifically, We assess the potential harmfulness of words by asking GPT-4o-mini (OpenAI (2024b), we use the 2024-07-18 version by default unless otherwise specified) questions like, “Do you think the word of {word} is especially frequent in malicious or prompt attack scenarios?” 1 Next, we perform a manual verification step, where we review the remaining words after LLM filtering to ensure that any words unrelated to prompt injection attacks are removed. Through these two steps, we have compiled a word list with the highest potential for malicious use.
@@
Corpus Generation. With a refined list of trigger words, we then generate sample sentences. We instruct the GPT-4o-mini to craft new sentences that must include a specified number of the selected trigger words 2. Specifically, we select the generated 113 trigger words and create three subsets with distinct difficulty levels, determined by the number of trigger words used in the generation of new sentences. Namely, three subsets separately containing 1, 2, and 3 trigger words are built with 113 benign sentences per subset. It is imperative that these sentences are contextually coherent, semantically meaningful, and do not contain any prompt injection instructions or malicious content. The goal is to represent benign usage of the trigger words in everyday language, ensuring that the sentences reflect natural and diverse linguistic patterns. To accomplish this, we design a carefully curated prompt (see Appendix. C.1) to guide the LLM in generating safe and natural sentences. We then implement a polish process to further ensure the safety of the generated samples. In this process, the generated sentences combined with the prompt in Appendix. C.2 are re-input into the LLM to identify any potential injection vulnerabilities. Following this, we conduct a manual review to confirm the safety of all sentences and report the error ratio of three subsets in Fig. 8. This multi-step refinement process guarantees that all generated sentences are harmless. The final output forms the proposed NotInject, which contains a total of 339 generated samples with 113 for one-word subset, 113 for two-word subset, and 113 for three-word subset. NotInject encompasses a diverse set of topics to enable a thorough evaluation, including common queries from daily life, technique queries (eg., programming, system), virtual creations, and multilingual queries (eg., Chinese, Russian, etc.). The detailed category distribution is presented in Table. 1.
Category one-word two-word three-word
Common Queries 58 49 19
Techniques 16 30 41
Virtual Creation 14 4 24
Multilingual 25 30 29
Table 1: Topic category distribution of NotInject.
footnotetext: 1, 2 Detailed prompts are in the Appendix. C
@@ 3.3 Evaluations on NotInject
Figure 6: Comparison on 3 dimensions of benign, malicious, and over-defense accuracy.
@@ 4 InjecGuard
In this section, we aim to develop a robust prompt guard model that excels in malicious accuracy, benign accuracy, and over-defense accuracy. We also want to emphasize that existing prompt guard models are typically trained in a closed-source manner and disclose their training data. Our goal is to train our model using only publicly available data, thereby promoting transparency and facilitating academic research. Here, we introduce InjecGuard, a prompt guard model that achieves state-of-the-art performance across various benchmarks. To build InjecGuard, we first curate a comprehensive collection of open-source datasets for training and apply data-centric augmentation techniques to address the long-tail problem within the dataset. More importantly, we propose a novel training strategy, MOF, which effectively mitigates over-defense issues without relying on specific over-defense datasets.
@@ 4.1 Data Collection and Augmentation
Similar to existing models, we first collect a wide range of prompt injection and benign corpus as the training data. Specifically, we select $20$ open-source datasets, including the benign datasets like Alpaca Taori et al. (2023), the prompt injection datasets like Safeguard-Injection Erdogan et al. (2024), and the datasets generated by existing injection attack, such as TaskTrack Abdelnabi et al. (2024)). However, upon analyzing these datasets, we identify a long-tail issue in these datasets: certain input formats frequently exploited in prompt injection attacks—such as CSV files—are underrepresented. To address this, we implement a data-centric augmentation procedure, generating additional data for these long-tail formats. Using GPT-4o-mini, we create some prompt injection samples in 17 formats, including Email, Document, Chat Conversation, JSON, Code, Markdown, HTML, URL, Base64, Table, XML, CSV, Config File, Log File, Image Link, Translation, and Website. After augmentation, our final training dataset for InjecGuard comprises $61,089$ benign samples and $15,666$ prompt injection samples, where $435$ samples are generated by our data-centric augmentation procedure. Detailed statistics of the training data and prompt for data-centric augmentation are provided in the Appendix. A and C.3, respectively.
@@ 4.2 Mitigating Over-defense for Free (MOF)
After training our model with the aforementioned data (Sec. 4.1) using standard supervised learning, we perform a “token-wised recheck” to identify biases in the trained model. This procedure takes every token in the tokenizer vocabulary and inputs each token individually into the trained model. Ideally, these tokens should all be considered “benign” since they are individual tokens without any intent of prompt injection attack. By doing this, we can identify any biased words that are incorrectly predicted as “attack” by the model. We consider the bias toward these tokens as the root cause of the over-defense issue in the model.
After identifying the biased tokens, we prompt GPT-4o-mini to generate benign data (see Appendix. C.1) using random combinations of these tokens (including one-token, two-token, and three-token settings, as mentioned in Sec. 3.2). We generate 1,000 benign samples using this method. Afterward, a LLM refinement process similar to the method mentioned in Sec. 3.2 is employed to ensure the harmlessness of generated data. Subsequently, we incorporate these generated data into the training data introduced in Sec. 4.1 to form the final training dataset. Using this dataset, we retrain our prompt guard model from scratch and get the final InjecGuard.

### v3

@@ Abstract
Prompt injection attacks pose a critical threat to large language models (LLMs), enabling goal hijacking and data leakage. Prompt guard models, though effective in defense, suffer from over-defense—falsely flagging benign inputs as malicious due to trigger word bias. To address this issue, we introduce NotInject, an evaluation dataset that systematically measures over-defense across various prompt guard models. NotInject contains 339 benign samples enriched with trigger words common in prompt injection attacks, enabling fine-grained evaluation. Our results show that state-of-the-art models suffer from over-defense issues, with accuracy dropping close to random guessing levels (60%). To mitigate this, we propose InjecGuard, a novel prompt guard model that incorporates a new training strategy, Mitigating Over-defense for Free (MOF), which significantly reduces the bias on trigger words. InjecGuard demonstrates state-of-the-art performance on diverse benchmarks including NotInject, surpassing the existing best model by 30.8%, offering a robust and open-source solution for detecting prompt injection attacks. The code and datasets are released at https://github.com/leolee99/InjecGuard.
@@
Hao Li1thanks: These authors contribute equally to this work., Xiaogeng Liu2 1Washington University in St. Louis, 2University of Wisconsin-Madison
@@ 1 Introduction
Prompt injection attacks Perez and Ribeiro (2022); Greshake et al. (2023); Liu et al. (2024) represent a serious and emerging threat to the security and integrity of large language models (LLMs) Brown et al. (2020). These attacks exploit the models’ reliance on natural language inputs by inserting malicious or manipulative prompts, leading to undesirable behaviors such as goal hijacking or sensitive data leakage. For instance, a well-known prompt injection technique involves instructing the LLM to “ignore previous instructions”, Branch et al. (2022); Harang (2023a); Perez and Ribeiro (2022); Willison (2022) which can override built-in safeguards and enable the execution of unauthorized actions.
@@
To address it, prompt guard models Meta (2024); ProtectAI.com (2024); Deepset (2024b); fmops (2024); LakeraAI (2024a) have recently been proposed as a promising solution. These models work by analyzing the semantic meaning of the input data to detect malicious intent before it reaches the LLM. Unlike LLMs, prompt guard models are lightweight and computationally efficient, as they do not require the high inference cost associated with LLMs. Additionally, these models operate independently, without the need for victim LLM’s responses, further reducing computation costs compared to approaches like LLM guardrails Inan et al. (2023). These attributes make prompt guard models adaptable to various environments, and an attractive solution in scenarios where speed and resource optimization are critical.
@@
Figure 2: Over-denfese issue in ProtectAIv2 ProtectAI.com (2024) 1, the current SotA prompt guard model.
@@
11footnotetext: https://huggingface.co/protectai/deberta-v3-base-prompt-injection-v2
@@ 2 Related Works
Prompt injection attacks. The concept of prompt injection attacks is first identified in research by Perez and Ribeiro (2022), revealing that LLMs could be misled by simple, crafted inputs, resulting in goal hijacking and prompt leakage. Several studies have been proposed Greshake et al. (2023); Wang et al. (2023); Pedro et al. (2023); Yan et al. (2023); Yu et al. (2023); Salem et al. (2023); Yip et al. (2024); Zhan et al. (2024); Liu et al. (2024); Pasquini et al. (2024); Shi et al. (2024), addressing various aspects of prompt injection attacks, such as handcrafted methods Toyer et al. (2023), automatic attack algorithms Liu et al. (2024), and benchmarks Liu et al. (2023); Debenedetti et al. (2024). Public discussions Harang (2023b); Willison (2022, 2023) have also underscored the risks of prompt injection attacks on commercial LLMs. Prompt Injection datasets are also introduced, such as PINT LakeraAI (2024b), Safeguard-Injection Erdogan et al. (2024), TaskTracker Abdelnabi et al. (2024), and BIPIA Yi et al. (2023), etc.
@@
Prompt guard models. Prompt guard models aim to detect malicious intent in inputs. These methods are computationally efficient and, unlike LLM guardrails Inan et al. (2023), do not require an additional round of victim LLM inference to generate a response. Several prompt guard models have been proposed recent including open-sourced Fmops fmops (2024), Deepset Deepset (2024b), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024), and commercial close-sourced LakeraGuard LakeraAI (2024a). However, these models are either trained on closed datasets or do not release their training details.
@@
Most of these prompt guard models fmops (2024); Deepset (2024b); Meta (2024); ProtectAI.com (2024) suffer from over-defense issues, where they rely on shortcuts triggered by certain keywords to make predictions, leading them to incorrectly categorize a benign input that contains the keywords as malicious. Moreover, all of these models are trained using closed-source data with undisclosed implementation or training details. In this paper, we introduce an over-defense dataset to systematically evaluate the over-defense issue, and propose a novel and completely open-source training approach that achieves SotA performance across diverse benchmarks, including the over-defense dataset.
Shortcut learning. Shortcut learning refers to the common phenomenon across various tasks Niven and Kao (2019); Wang et al. (2021); Manjunatha et al. (2019), where machine learning models Geirhos et al. (2020); Alshemali and Kalita (2020) develop spurious correlations between input features and target labels. To address this challenge, a range of mitigation strategies have been proposed, including regularization Wang et al. (2021), domain adaptation Liu and Avci (2019), data augmentation Wu et al. (2022), multi-task learning Du et al. (2021). However, recent studies suggest that these techniques may sometimes undermine model performance Khani and Liang (2021), particularly in in-domain distribution (IID) scenarios Mendelson and Belinkov (2021). Unlike many previous tasks Niven and Kao (2019); Wang et al. (2021); Manjunatha et al. (2019); Lin et al. (2024), injection attack detection is both a semantic- and pattern-based task, making it more vulnerable to spurious correlations compared to semantic-only tasks. Effectively mitigating these correlations without sacrificing performance is a significant challenge for this task.
@@ 3 Over-defense Dataset: NotInject
Figure 4: The pipeline for constructing NotInject dataset
@@ 3.1 The Over-defense Issue
While prompt guards offer several advantages, such as low overhead as previously mentioned, we have identified a critical limitation: they exhibit severe over-defense issues, even in some advanced models Meta (2024); ProtectAI.com (2024). Specifically, these models learn a shortcut from certain trigger words, like “ignore”, directly to the final prediction. As shown in Fig. 3 with red lines, ProtectAIv2 ProtectAI.com (2024), one of the SotA prompt guard models on PINT benchmark, distributes excessive and unbalanced attention to the word “ignore”, leading it to incorrectly categorize a benign input as malicious. To further illustrate, we input a benign sentence containing the word “ignore” into two advanced prompt guard models, ProtectAIv2 ProtectAI.com (2024) and PromptGuard Meta (2024), with the results separately shown in Fig. 2 and Fig. 8. Both models incorrectly classifies the benign instruction containing the word “ignore” as malicious.
@@ 3.2 Construction of NotInject
Trigger Words Refinement. Recognizing that automated methods may include irrelevant or common words not indicative of prompt injection attempts, we proceed with a word rectification process. We first employ LLM to automatically filter the list of candidate words to remove any that may not be pertinent. Specifically, We assess the potential harmfulness of words by asking GPT-4o-mini (OpenAI (2024b), we use the 2024-07-18 version by default unless otherwise specified) questions like, “Do you think the word of {word} is especially frequent in malicious or prompt attack scenarios?” Next, we perform a manual verification step, where we review the remaining words after LLM filtering to ensure that any words unrelated to prompt injection attacks are removed. Specifically, we employ three human evaluators with security expertise to assist in the refinement process. Each evaluator independently scores the frequency of each word based on the agreement shown in Fig. 6. The average score for each word is then calculated, and words with an average score above 3 are identified as trigger words. Through these two steps, we have compiled a word list with the highest potential for malicious use.
@@
Corpus Generation. With a refined list of trigger words, we then generate sample sentences. We instruct the GPT-4o-mini to craft new sentences that must include a specified number of the selected trigger words. Specifically, we select the generated 113 trigger words and create three subsets with distinct difficulty levels, determined by the number of trigger words used in the generation of new sentences. Namely, three subsets separately containing 1, 2, and 3 trigger words are built with 113 benign sentences per subset. It is imperative that these sentences are contextually coherent, semantically meaningful, and do not contain any prompt injection instructions or malicious content. The goal is to represent benign usage of the trigger words in everyday language, ensuring that the sentences reflect natural and diverse linguistic patterns. To accomplish this, we design a carefully curated prompt (see Appendix. D.1) to guide the LLM in generating safe and natural sentences. We then implement a polish process to further ensure the safety of the generated samples. In this process, the generated sentences combined with the prompt in Appendix. D.2 are re-input into the LLM to identify any potential injection vulnerabilities. Following this, we conduct a manual review to confirm the safety of all sentences and report the error ratio of three subsets in Fig. 10. This multi-step refinement process guarantees that all generated sentences are harmless. The final output forms the proposed NotInject, which contains a total of 339 generated samples with 113 for one-word subset, 113 for two-word subset, and 113 for three-word subset. NotInject encompasses a diverse set of topics to enable a thorough evaluation, including common queries from daily life, technique queries (eg., programming, system), virtual creations, and multilingual queries (eg., Chinese, Russian). The detailed category distribution is presented in Tab. 8.
@@ 3.3 Evaluations on NotInject
Figure 5: Comparison of benign, malicious, and over-defense accuracy across various prompt guard solutions. InjecGuard significantly outperforms all prior solutions. Notably, the open-source models (Deepset, Fmops, PromptGuard, ProtectAIv2) exhibit significant over-defense issues, with over-defense accuracy under 60%, where 50% represents random guessing. In addition, although LakeraGuard and LlamaGuard3 demonstrate strong over-defense performance, their effectiveness is still limited by suboptimal malicious accuracy.
@@ 4 InjecGuard
We introduce InjecGuard, a prompt guard model with better performance in terms of malicious accuracy, benign accuracy, and over-defense accuracy. First, we will introduce data collection and augmentation, followed by our novel training strategy.
@@ 4.1 Data Collection and Augmentation
Almost all existing injection guard solutions (e.g., FMops fmops (2024), PromptGuard Meta (2024), ProtectAIv2 ProtectAI.com (2024)) have only open-sourced their models, but the datasets and implementation details used for training remain closed-source. This means that there is no off-the-shelf public dataset available for training injection guard models. Therefore, the first step is to collect a wide range of prompt injection and benign corpus as a basic training dataset. Specifically, we select $20$ open-source datasets, including the benign datasets like Alpaca Taori et al. (2023), the prompt injection datasets like Safeguard-Injection Erdogan et al. (2024), and the datasets generated by existing injection attack, such as TaskTrack Abdelnabi et al. (2024)). However, upon analyzing these datasets, we identify a long-tail issue in these datasets: certain input formats frequently exploited in prompt injection attacks—such as CSV files—are underrepresented. To address this, we implement a data-centric augmentation procedure, generating additional data for these long-tail formats. Using GPT-4o-mini, we create some prompt injection samples in 17 formats, including Email, Document, Chat Conversation, JSON, Code, Markdown, HTML, URL, Base64, Table, XML, CSV, Config File, Log File, Image Link, Translation, and Website. After augmentation, our final training dataset for InjecGuard comprises $61,089$ benign samples and $15,666$ prompt injection samples, where $435$ samples are generated by our data-centric augmentation procedure. Detailed statistics of the training data and prompt for data-centric augmentation are provided in the Appendix. A and D.3.
@@ 4.2 Mitigating Over-defense for Free (MOF)
After training our model with the aforementioned data (Sec. 4.1) using standard supervised learning, we perform a “token-wised recheck” to identify biases in the trained model. This procedure takes every token in the tokenizer vocabulary and inputs each token individually into the trained model. Ideally, these tokens should all be considered “benign” since they are individual tokens without any intent of prompt injection attack. By doing this, we can identify any biased words that are incorrectly predicted as “attack” by the model trained on basic datasets in Sec. 4.1. We consider the bias toward these tokens as the root cause of the over-defense issue in the model.
After identifying the biased tokens, we prompt GPT-4o-mini to generate benign data (see Appendix. D.1) using random combinations of these tokens (including one-, two-, and three-token settings, as mentioned in Sec. 3.2). We generate 1,000 benign samples (the scale exploration see Appendix. B.2) using this method. Afterward, a LLM refinement process similar to the method mentioned in Sec. 3.2 is employed to ensure the toxicity of generated data. Subsequently, we incorporate these generated data into the training data introduced in Sec. 4.1 to form the final training dataset. Using this dataset, we retrain our prompt guard model from scratch and get the final InjecGuard.